OpenStudy (anonymous):

Simplify the expression. (1 - cot x) / tan x - 1

OpenStudy (anonymous):

substitute the value of cot x and tan x i hope u will get the ans

OpenStudy (anonymous):

What?

OpenStudy (jdoe0001):

well, keeping in mind that cot= cos/sin, and tan = sin/cos let's use that atop and below $$ \cfrac{1-cot(\theta)}{tan(\theta)-1} \implies \cfrac{ 1-\frac{cos(\theta)}{sin(\theta)} } { \frac{sin(\theta)}{cos(\theta)} } $$

OpenStudy (jdoe0001):

hemm I missed something :/

OpenStudy (anonymous):

-2

OpenStudy (jdoe0001):

$$ \cfrac{1-cot(\theta)}{tan(\theta)-1} \implies \cfrac{ 1-\frac{cos(\theta)}{sin(\theta)} } { \frac{sin(\theta)}{cos(\theta)}-1 } $$

OpenStudy (jdoe0001):

so, there, add both, atop and bottom, what do you get?

OpenStudy (anonymous):

Add what?

OpenStudy (jdoe0001):

the fractions :/

OpenStudy (anonymous):

How?

OpenStudy (jdoe0001):

$$ \cfrac{1-cot(\theta)}{tan(\theta)-1} \implies \cfrac{ 1-\frac{cos(\theta)}{sin(\theta)} } { \frac{sin(\theta)}{cos(\theta)}-1 } \implies \large { \color{red}{ \cfrac{ \frac{1}{1}-\frac{cos(\theta)}{sin(\theta)} } { \frac{sin(\theta)}{cos(\theta)}-\frac{1}{1} }} } $$

OpenStudy (jdoe0001):

let's try first the numerator

OpenStudy (anonymous):

(1 - cos(x))/1 - sin(X) ?

OpenStudy (jdoe0001):

\(\frac{1}{1}-\frac{cos(\theta)}{sin(\theta)}\) what would be your LCD there?

OpenStudy (anonymous):

sin (x)

OpenStudy (jdoe0001):

right, so let's use that, to get $$ \frac{1}{1}-\frac{cos(\theta)}{sin(\theta)} \implies \frac{sin(\theta)-cos(\theta)}{sin(\theta)} $$

OpenStudy (anonymous):

Ok. then the denominator would just be cos (x) - sin (x) / cos (x)

OpenStudy (jdoe0001):

right

OpenStudy (jdoe0001):

well, one sec

OpenStudy (jdoe0001):

the denominator will be sin-cos/cos

OpenStudy (anonymous):

right. sorry.

OpenStudy (anonymous):

so sin-cos/sin / (sin-cos/cos)?

OpenStudy (jdoe0001):

$$\large { \cfrac{ \frac{sin(\theta)-cos(\theta)}{sin(\theta)} }{ \frac{sin(\theta)-cos(\theta)}{cos(\theta)} } \implies \frac{sin(\theta)-cos(\theta)}{sin(\theta)} \times \frac{cos(\theta)}{sin(\theta)-cos(\theta)} } $$

OpenStudy (jdoe0001):

so, there are 2 guys to cancel out, and you're left with just a function :)

OpenStudy (anonymous):

cot x

OpenStudy (anonymous):

Ok. Thank you!! :)

OpenStudy (jdoe0001):

yw